OpenStudy (anonymous):

The radius of a puddle is claimed to be 12.0 inches, correct to within 0.01 inch. Use linear approximation to estimate the resulting error, measured in square inches, in the area of the puddle. ± 0.24π ± 0.024π ± 0.06π ± 0.0004166

OpenStudy (dan815):

uhh

OpenStudy (dan815):

0.01/12 = relative error

OpenStudy (dan815):

A=pi*r^2

OpenStudy (dan815):

hmm

OpenStudy (anonymous):

I used the formula for the area of a circle A=pi*r^2 dA=2pi*r to get the linear approx. at x=12 A=f(12)+f'(12)(x-12) =144pi+24pi(x-12) 24pi(x-6) skipped some simplifying steps

OpenStudy (anonymous):

I put the answers up

OpenStudy (anonymous):

I keep getting 0.0164493

OpenStudy (anonymous):

Oh wait nvm! @dan815 I got it

OpenStudy (anonymous):

You find the area of the puddle for 12 inches then you pick either 12.01 or 11.99 as another radius and subtract 12^2*pi from it and you get +/- 0.24pi

OpenStudy (anonymous):

12=144pi 12.01=12.01^2*pi 12.01^2*pi-144pi=0.24pi

OpenStudy (dan815):

ok

OpenStudy (dan815):

(0.01/12)*2*144pi=0.24 pi

OpenStudy (dan815):